Cowling’s rule. Cowling’s rule is another method for
determining the dosage of a drug to prescribe to a child.
For this rule, the formula
d=D(a+1) divided by 24
gives the child’s dosage d, where D is the adult dosage
and a is the age of the child in years. If the adult dosage
of a drug is 600 milligrams and a doctor uses this
formula to determine that a child’s dosage is
200 milligrams, then how old is the child?

Answer by stanbon(75887)

the formula d=D(a+1) divided by 24 gives the child’s
dosage d, where D is the adult dosage and a is the
age of the child in years.
---------------------------------------------------------
If the adult dosage of a drug is 600 milligrams and
a doctor uses this formula to determine that a child’s
dosage is 200 milligrams, then how old is the child?
-----------------------------
d=D(a+1)/24
-----
200 = 600(a+1)/24
-----
a+1 = 8
------
a = 7 yrs old
==========
Cheers,
Stan H.